OpenStudy (anonymous):

The mean percentage of a population of people eating out at least once a week is 57%, with a standard deviation of 3.5%. Assume that a sample size of 40 people was surveyed from the population an infinite number of times. In which interval will 95% of the sample means occur?

OpenStudy (kropot72):

Taking samples of 40 people an infinite number of times results in the distribution of the sample means closely approaching a Normal distribution with the following parameters: Mean = population mean Standard deviation = population standard deviation/square root of the sample size. The interval in which 95% of thesample means occur is as follows: \[(57-1.96\times \frac{3.5}{\sqrt{40}}\ ,\ 57+1.96\times \frac{3.5}{\sqrt{40}})\]

OpenStudy (anonymous):

im still confused, like where do i go from there?

OpenStudy (anonymous):

@tcarroll010

OpenStudy (kropot72):

@pascuzzoa Just calculate the two values inside the brackets.

OpenStudy (anonymous):

@pascuzzoa on one side of the comma is an equation on the other is a seperate equation they will give you seperate values think of it as a minimum and a maximum this is the answer between 55.89% and 58.11%

OpenStudy (anonymous):

@kropot72

OpenStudy (kropot72):

@levils Basically, you are correct.